OpenStudy (cruffo):

The greatest integer function is defined by \[ \lfloor x \rfloor = \text{the largest integer that is less than or equal to x}\]Let \[f(x) = \lfloor x \rfloor + \lfloor -x \rfloor\] Show that \[\lim_{x \to 2}\; f(x)\] exists but is not equal to f(2).

OpenStudy (cruffo):

Repost: The greatest integer function is defined by \[ \lfloor x \rfloor = \text{the largest integer that is less than or equal to x}\]Let \[f(x) = \lfloor x \rfloor + \lfloor -x \rfloor\] Show that \[\lim_{x \to 2}\; f(x)\] exists but is not equal to f(2).

OpenStudy (watchmath):

Note that \(\lfloor x\rfloor +\lfloor -x\rfloor\leq x+(-x+1)=1\)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2^+}f(x)=[2^+]+[-2^+]=2-3=-1\]\[\lim_{x \rightarrow 2^-}f(x)=[2^-]+[-2^-]=1-2=-1\] f(2)=0

OpenStudy (watchmath):

dipankar is right :)

OpenStudy (watchmath):

we don't need to do in general since we are only interested for x near 2.

OpenStudy (cruffo):

That was a quick answer!